What volume of NO gas at STP can be produced from 7.0 moles of nitrogen dioxide (NO₂) according to the following equation? 3NO₂ + H₂O → 2HNO₃ + NO *
1
Expert's answer
2020-04-27T01:53:31-0400

The balanced equation for considered reaction can be written as follows

"3NO_2 + H_2O \rightarrow 2HNO_3 + NO"


"\frac{\nu (NO_2)}{n(NO_2)} = \frac{V(NO)}{V_M \cdot n(NO)}"

where "\nu" - amount, mol; "V" - volume, L; "V_M" - molar volume of a gas, "V_M=22.4 \; L/mol" ; "n" - number of molecules in the balanced equation.

"V(NO)=\frac{\nu (NO_2) \cdot V_M \cdot n(NO)}{n(NO_2)}"

"V(NO)=\frac{7.0 \cdot 22.4 \cdot 1}{3} \approx 52.3 \; L"


Answer: 52.3 L.